OpenStudy (anonymous):

The length of a rectangle is 3 centimeters more than 3 times the width. If the rectangle's perimeter is 46 centimeters, what are the rectangle's dimensions?

OpenStudy (anonymous):

P = 2L + 2W = 2(3W + 3) + 2w = 8W + 6 = 46 8W = 40 W = 5 L = (46 - 10)/2 = 18

OpenStudy (anonymous):

5 by 18?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thnx, bro

OpenStudy (anonymous):

You're welcome! thx for the recognition!

OpenStudy (anonymous):

np